40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 13850     Entry time: Wed May 16 21:47:17 2018
Author: Kevin 
Type: Update 
Category: PEM 
Subject: Seismometer Noise Spectra 

Earlier today Kira and I reconnected the EX seismometer. I just took some spectra of all three seismometers, shown in the attachments, to compare with past data and to do a rough check of the calibration.

This elog has a spectra from 2010 (GUR1 is now EY) and this elog has one for BS at lower frequencies from 2017. Note that the EX seismometers now have strong peaks that are not at 60 Hz harmonics. Other than these peaks, these old spectra roughly match up with the ones taken today, so the callibration is still roughly the same. I couldn't find any old data for EX (GUR2) though so I don't know for sure that these peaks weren't there before.


gautam 20180517 0930: In 2017, Gur2 (now EX) looked like this. Still peaky, but the peaks seem shifted in frequency. Steve also informed me that the Gur1 and Gur2 cables were swapped n times, so perhaps we shouldn't read too much into that.

Attachment 1: BS_vel.pdf  221 kB  Uploaded Wed May 16 22:50:02 2018  | Hide | Hide all
BS_vel.pdf
Attachment 2: EX_vel.pdf  221 kB  Uploaded Wed May 16 22:50:39 2018  | Hide | Hide all
EX_vel.pdf
Attachment 3: EY_vel.pdf  221 kB  Uploaded Wed May 16 22:50:54 2018  | Hide | Hide all
EY_vel.pdf
ELOG V3.1.3-